Main INDEX
Monthly INDEX
PREV
NEXT
Make New Entry,
Make Followup Entry
User name E.Chudakov
Log entry time 18:23:42 on November 21, 2008
Entry number 250215
keyword=Experiment summary for Nov 20(swing) - 21(day)
Summary for Nov 20 (swing) - Nov 21 (day)
Relatively smooth data taking
Beam in Use
shift |
ABU |
BANU |
BNA |
2008/11/20 swing | 5.40 |
0.78 | 1.83 |
2008/11/21 owl | 6.51 |
0.53 | 0.96 |
2008/11/21 day | 4.31 |
0.98 | 2.70 |
total | 16.22 |
2.29 | 2.79 |
Main events/issues:
- The left Q1 gave no trouble
- The dehumidifier's compressor started burning oil again. The VESDA
smoke alert came back. At 8:30 on Nov 21 an access was taken (45
min) and
the compressor was turned off till next Tuesday.
- 1/2 wave plane removed at about 9:10.
- RICH is in the DAQ data stream, giving no dead time trouble.
Plans for the next few days:
- Running with minimal changes;
- Monday: Compton test may be scheduled providing their readiness.
- Tuesday: target cell change and a bunch of lesser things, to be
coordinated.
A copy of this log entry has been emailed to: transversity@jlab.org